hugo is eight years older than gabby. in five years, hugo will be twice as old as gabby.how old is each now?
I know the ages are 11 and 3. I just can't get can't show work.

Answer by scott8148(6628)

"hugo is eight years older than gabby" ___ h = g + 8

"in five years, hugo will be twice as old as gabby" ___ h + 5 = 2(g + 5)

looks like substitution might be easiest